Use elimination to solve the system. (Simplify your answer completely.)
3x − 5y = 3
7x + 3y = 51

Answer by Boreal(15235)

3x − 5y = 3
7x + 3y = 51
multiply top by 3 and bottom by 5 and add them
9x-15y=9
35x+15y=255
44x=264
x=6
go to first
18-5y=3
-5y==15
y=3
(6, 3) ANSWER
check in second
42+9=51 checks